# ControlDreamer: Stylized 3D Generation with Multi-View ControlNet

### v2

@@ 3.1.2 Style Stage
Our second stage is motivated by the coarse-to-fine optimization of [21], which is utilized for generating high-resolution meshes. In their process, the textured mesh is optimized with DMTet algorithm [38], which uses a deformable tetrahedral grid and employs a differentiable rasterizer [19, 25] for efficient rendering. However, unlike the goal of [21], which is centered on high-resolution mesh refinement, our approach leverages the DMTet algorithm for generating style on the previously generated 3D model.
@@
In each iteration of optimization, we employ pre-trained MiDAS [31] to obtain predicted depth maps, leveraging the identical depth estimator that is used for training our MV-ControlNet in Sec. 3.2. Subsequently, we optimize mesh with the gradient computed by MV-ControlNet as shown in the following equation:
@@
$\nabla_{\theta}\mathcal{L}_{\text{SDS}-\text{C}}\triangleq\mathbb{E}_{t,\epsilon}\left[w(t)\left(\epsilon_{\phi}(z_{t};\mathcal{D}(x),y_{style},c,t)-\epsilon\right)\frac{\partial z}{\partial x}\frac{\partial x}{\partial\theta}\right],$ (2)
@@
where $y_{style}$ is the style text prompt, which is differently used with $y_{geo}$. $\mathcal{D}$ is our MV-ControlNet, which we later explain in Sec. 3.2. This gradient is coupled with gradients of regularizations to meticulously refine both geometry and texture. In contrast to recent methods [21, 5, 42] that necessitate narrow $[t_{\text{min}},t_{\text{max}}]$ during the refinement phase, our ControlNet allows a large range of $[t_{\text{min}},t_{\text{max}}]$ with scheduled timestep annealing. This ensures that depth information remains intact, safeguarding original geometry throughout the style optimization, even with significant variations in timesteps.
Similar to the iterative dataset update concept introduced in [11], we discovered that iteratively updating mesh with MiDAS-guided estimated depth maps effectively edits the geometry and texture during a single-stage training of the given source 3D models toward the desired target 3D models.
3.1.3 Regularizations
We use the following two regularizations along with score distillation sampling. First, we use normal map consistency loss [21], which ensures the mesh’s normal map remains consistent. Second, we use Laplacian smoothing regularization [42], which eliminates pixelated artifacts in DMTet. These allow us to generate photorealistic 3D models.
3.1.4 Annealing diffusion timesteps
@@
During optimization, we linearly anneal the timestep range of $t$. Utilizing MV-ControlNet, we have discovered that textures can be effectively modified while preserving the initial geometry, even with significant variations in text descriptions. Furthermore, this effectiveness persists in scenarios where $t$ is sampled from $t\sim U(0.02,0.98)$. We continued sampling $t$ from this range for the initial $4/5$ of the total iterations, akin to the geometry stage. Subsequently, for the remainder of the iterations, $t$ is sampled from $t\sim U(0.02,0.5)$. MV-ControlNet allows us to deviate from traditional methods that rely on smaller timesteps for mesh refinement.
Figure 4: Additional qualitative results. Our generation process involves first generating a coarse-grained geometry, and then generating a fine-grained stylized 3D model with a style prompt.
3.1.5 Editing the geometry and texture
ProlificDreamer [42] comprises two refinement stages: geometric refinement followed by texture refinement. Furthermore, each training iteration requires not only the computation of the SDS loss but also the fine-tuning of the diffusion model, resulting in a computationally expensive generation process.
@@ 3.2 Multi-View ControlNet
As elaborated in the previous section, our pipeline generates style in the second stage by refining the DMTet, which is converted from the NeRF. We note the limitations of MVDream in refining DMTet, stemming from a lack of understanding of the geometry. To address this, we introduce the Multi-View (MV) ControlNet, crucial for generating fine-grained style while comprehending the coarse-grained mesh’s geometry.
@@ 3.2.1 Generating Dataset
To create a depth-aware multi-view model, we use the curated 100K high-quality texts from OpenShape [22], filtered with GPT-4 [27] for quality assurance. These texts are first refined using BLIP [20] and Azure cognition services for initial 2D image captions, and then enriched by GPT-4, filtering out uninformative content. We utilize MVDream-generated multi-view images as training data, matching MVDream’s camera parameters to maintain its prior knowledge. This involves generating four orthogonal views for each text, with uniformly distributed azimuth angles and elevation. Specifically, we randomly sample the camera parameters in a range of [0.9, 1.1] for distance, [15, 60] for fov, and [0, 30] for elevation. Then, we generate multi-view depth maps from the generated images using MiDAS [31]. The process, completed in 1.5 days on four A40 GPUs, utilizes texts with explicit ‘3D’ mentions or appended them with a ‘3D asset’ postfix.
@@ 3.2.2 Training MV-ControlNet
MVDream [39] expands the text-to-image capabilities of Stable Diffusion [33] to enable text-to-multi-view generation. The model processes inputs of noisy latents $z_{t}\in\mathbb{R}^{F\times H\times W\times C}$, camera parameters $c\in\mathbb{R}^{F\times 16}$, and text prompts. MVDream differentiates itself by utilizing a 3D self-attention mechanism, an extension of 2D self-attention, to model multi-view consistency and capture the interrelations among pixels across $F$ views. Note that additional operations such as cross-attentions and convolutions operate on a frame-wise basis.
Our goal is to train a depth encoder that incorporates multi-view depth conditions into MVDream, leveraging the model’s multi-view consistency and generation quality. Following the ControlNet [43] approach, which involves training an additional encoder atop a frozen diffusion model, we introduce multi-view depths as additional inputs to the pre-trained MVDream. Specifically, a trainable copy of MVDream’s U-Net [34] encoder, including 3D attentions, is tasked with encoding depth and extracting multi-resolution features, which are then integrated into the U-Net decoder’s skip connections using zero-initialized convolutions. We train this encoder with the epsilon prediction objectives [13], formalized as follows:
@@
where $\mathcal{E}$ denotes the VAE encoder of Stable Diffusion [33], $\mathcal{D}$ is the trainable depth encoder, $\epsilon_{\phi}$ is the frozen MVDream’s U-Net, $y$ is the text prompt, and $c$ is the camera parameter.
Figure 5: We compare the depth-aware MV-ControlNet with [12] approach using MVDream, as well as MV-ControlNet variants trained under the edges and normals conditions. The left side shows source images along with their conditions. The depth-conditioned multi-view images consistently display the most visually appealing results.
Figure 6: The left column displays an array of initial 3D models. From each initials, multiple styles are applied to create 3D models. Utilizing a two-stage pipeline, this process generates customized 3D models with a variety of geometries, including samurai, Mr. Bean, stone bust, and a Messi figure, and styles including Iron Man, Captain America, and Superman. Our ControlDreamer demonstrates visually appealing results for diverse, and challenging concepts.
@@ 4 Experiments
In our experiments, we carry out tests using user-customized prompts, distinct from the $100K$ text corpus used for training our MV-ControlNet. For both qualitative and quantitative assessments, we have created a benchmark dataset comprising a total of 30 pairs of geometry and style prompts specifically for 3D editing experiments. These pairs span across five distinct categories: Animals, Characters, Foods, General, and Objects. The main qualitative samples of our pipeline are presented in Fig. 6.
@@ 4.1 Qualitative Comparison of Multi-View Image Generation
To evaluate our MV-ControlNet’s depth-aware multi-view image generation capabilities, we conduct comparisons with several baselines as shown in Fig. 5. The baselines include MVDream using the prompt-to-prompt (P2P) [12] method and MV-ControlNet variants trained under the canny edge and normal conditions. P2P is an image editing method that aligns source and target images’ geometries by injecting attention maps into diffusion models. We use injection ratios set at 0.4 for cross-attention and 0.8 for self-attention, ensuring strong geometric alignment. However, the applicability of P2P is somewhat constrained due to the need for equal token lengths in both source and target texts, often resulting in less aligned text-guided editing outcomes. Conversely, as the proposed MV-ControlNet can ensure source geometry with depth conditions, it is not restricted by the length of the text, enabling the user’s application of more creative and diverse prompts in the style stage. Additional comparative results are displayed in Fig. 15 in the Appendix.
Regarding MV-ControlNet variants, we observe that canny edge maps result in unrealistic and 3D-inconsistent images, whereas normal maps provide 3D consistency but show degraded text alignment. Conversely, using depth conditions is the most effective, leading to visually appealing and well-aligned results. This success motivated us to choose depth-conditioned MV-ControlNet as our model. Additional experimental details and results are available in the appendix.
Figure 7: Human preference study on editability and text alignment in five prompt domains of our benchmark. ControlDreamer consistently ranks highest in user preferences across all domains.
Figure 8: The left plot illustrates the 3D model generation results at 45 degrees and 135 degrees under various input conditions. On the right, (a) demonstrates the refinement using Magic3D, and (b) highlights the results achieved with our ControlDreamer.
Figure 9: Comparison with previous text-to-3D pipelines. Hulk’s geometry from Fig. 1 is styled with Ironman and Spiderman’s styles respectively. Magic3D and ProlificDreamer often show texture artifacts due to misalignment between style prompts and the initial geometry. Fantasia3D and MVDream tend to result in oversaturation or residual appearance traits.
@@ 4.2 Qualitative Comparison of 3D Generation
We compare against several baselines, namely Magic3D [21], Fantasia3D [5], and ProlificDreamer [42]. All these methods, including our ControlDreamer, refine textured meshes, employing the DMTet [38] algorithm for mesh generation.
The results of our stylization experiments using these baseline models are illustrated in Fig. 9. With Magic3D and ProlificDreamer, we observe some instances of failure, characterized by only partial texture changes in the geometry. This issue arises from a misalignment between the prior knowledge embedded in Stable Diffusion for the style prompt (e.g., the Spiderman is slim) and the geometry from the previous stage. Our MV-ControlNet addresses this misalignment by understanding the depth information of the geometry, thereby enhancing alignment and coherence. While Fantasia3D and MVDream successfully stylize the overall geometry, they tend to produce results with either oversaturation or residual appearance traits from the original geometry prompt, such as persistent green skin color.
@@
4.3 Quantitative Comparison of 3D Generation
4.3.1 Directional CLIP Similarity
To quantitatively compare the alignment between prompt modifications and image changes in ControlDreamer and MVDream, we evaluate the directional CLIP similarity [9, 16] between first and second stage 3D models on our benchmark dataset. This score measures how well the editing directions of rendered images align with the text directions within the CLIP space. We use texts with three view-oriented postfixes: ’a front view’, ’a side view’, and ’a back view’, utilizing four images from each camera view (0°, 90°, 180°, 270° azimuth angles) for the generated 3D models.
In Table 2, averaged similarities and standard deviations are presented for each domain, across all six prompt candidates and directions. We employ the OpenCLIP ViT-L/14 model [7], trained on the LAION-2B dataset, which is different from the text encoder used in ControlDreamer. Our method demonstrates superior performance compared to other baselines, excelling in generating text-aligned 3D models across all domains in our benchmark. Detailed explanations of the used prompts are presented in Table. A. 3 in Appendix.
@@ 4.3.2 Human Evaluations
We assess preference scores using Amazon Mechanical Turk (AMT) across five domains and present them in Fig. 7. As a result, the proposed method consistently achieves the highest preference scores in all domains, demonstrating its effectiveness. This assessment allows for a comparative ranking of MVDream, Fantasia3D, Magic3D, and ProlificDreamer. We assess each methodology using 30 3D models from each domain, with 200 human evaluators participating. Each evaluator encounters three random samples and repeats the evaluation task three times, resulting in about 20 evaluations per domain. For additional details, the templates for the human evaluation are included in the Appendix.
@@ 4.4.1 Variants of MV-ControlNet
In Fig. 8, we compare the 3D stylization performance of MV-ControlNet variants trained under canny edge, normal, and depth conditions. With canny edge inputs, significant artifacts appear in both RGB and normal maps. When using normal maps, the RGB colors in the 3D models are less photorealistic. However, depth maps consistently produce 3D models without artifacts in both color and normal maps, showcasing their superiority.

### v3

@@ 3.1.2 Style Stage
Our second stage is motivated by the coarse-to-fine optimization of [Lin et al.(2023)Lin, Gao, Tang, Takikawa, Zeng, Huang, Kreis, Fidler, Liu, and Lin], which is utilized for generating high-resolution meshes. In this process, they optimize the textured mesh using DMTet [Shen et al.(2021)Shen, Gao, Yin, Liu, and Fidler], which uses a deformable tetrahedral grid and employs a differentiable rasterizer [Laine et al.(2020)Laine, Hellsten, Karras, Seol, Lehtinen, and Aila, Munkberg et al.(2022)Munkberg, Hasselgren, Shen, Gao, Chen, Evans, Müller, and Fidler] for efficient rendering. However, unlike the goal of [Lin et al.(2023)Lin, Gao, Tang, Takikawa, Zeng, Huang, Kreis, Fidler, Liu, and Lin], which is centered on high-resolution mesh refinement, our approach leverages the DMTet algorithm for generating style on the previously generated geometry.
@@
In each iteration of optimization, we employ pre-trained MiDAS [Ranftl et al.(2022)Ranftl, Lasinger, Hafner, Schindler, and Koltun] to obtain predicted depth maps, using the same depth estimator that is used for training our MV-ControlNet in Sec. 3.2. Subsequently, we optimize the mesh with the gradient computed by MV-ControlNet as shown in the following equation:
@@
$\nabla_{\theta}\mathcal{L}_{\text{SDS}-\text{C}}\triangleq\mathbb{E}_{t,\epsilon}\left[w(t)\left(\hat{\epsilon}_{\phi}(z_{t};\mathcal{D}(x),y_{\text{style}},c,t)-\epsilon\right)\frac{\partial z}{\partial x}\frac{\partial x}{\partial\theta}\right],$ (2)
@@
where $y_{\text{style}}$ is the style text prompt, which is used differently from $y_{\text{geo}}$. $\mathcal{D}$ is our MV-ControlNet, which we later explain in Sec. 3.2. This gradient is coupled with regularizations to refine both geometry and texture meticulously. In contrast to recent methods [Lin et al.(2023)Lin, Gao, Tang, Takikawa, Zeng, Huang, Kreis, Fidler, Liu, and Lin, Chen et al.(2023b)Chen, Chen, Jiao, and Jia, Wang et al.(2023)Wang, Lu, Wang, Bao, Li, Su, and Zhu] that necessitate narrow $[t_{\text{min}},t_{\text{max}}]$ during the refinement phase, our ControlNet allows a large range of $[t_{\text{min}},t_{\text{max}}]$ with scheduled timestep annealing. This approach ensures the depth information remains intact maintaining original geometry throughout the style optimization, even with significant timestep variations.
3.1.3 Training Details
@@
We use the normal map consistency loss [Lin et al.(2023)Lin, Gao, Tang, Takikawa, Zeng, Huang, Kreis, Fidler, Liu, and Lin] and the Laplacian smoothing regularization loss [Wang et al.(2023)Wang, Lu, Wang, Bao, Li, Su, and Zhu] to eliminate pixelated artifacts in DMTet. These allow us to generate photorealistic 3D models. Utilizing MV-ControlNet, textures are effectively modified while preserving the initial geometry, even with significant variations in text descriptions. Furthermore, this effectiveness persists in scenarios where $t$ is sampled from $t\sim U(0.02,0.98)$. In the style stage, we sample $t$ from this range for the initial $4/5$ of the total iterations, akin to the geometry stage. Subsequently, for the remaining iterations, $t$ is sampled from $t\sim U(0.02,0.5)$. These techniques differentiate MV-ControlNet from previous methods [Lin et al.(2023)Lin, Gao, Tang, Takikawa, Zeng, Huang, Kreis, Fidler, Liu, and Lin, Chen et al.(2023b)Chen, Chen, Jiao, and Jia] that rely on smaller timesteps for mesh refinement.
3.1.4 Theoretical Discussion
The key to the success of score distillation in text-to-3D generation lies in the guidance mechanism [Ho and Salimans(2021)]. The guided score of Eq. 1 can be written as the following equation:
$\hat{\epsilon}_{\phi}(z_{t};y_{\text{geo}})$ $=s\cdot(\epsilon_{\phi}(z_{t};y_{\text{geo}})-\epsilon_{\phi}(z_{t}))+\epsilon_{\phi}(z_{t})$ (3)
$\propto\nabla_{z_{t}}(s\cdot(\text{log}p(z_{t}|y_{\text{geo}})-\text{log}p(z_{t}))+\text{log}p(z_{t}))$ (4)
$=\nabla_{z_{t}}(s\cdot\text{log}p(y_{\text{geo}}|z_{t})+\text{log}p(z_{t})),$ (5)
due to Bayes’ rule. Here, $s$ represents the guidance scale and $c$ and $t$ are omitted for brevity. The equation indicates strong guidance encourages higher text alignment $p(y_{\text{geo}}|z_{t})$.
The effect of integrating our MV-ControlNet into score distillation can be demonstrated similarly. The guided score in Eq. 2 can be expressed as follows:
$\hat{\epsilon}_{\phi}(z_{t};\mathcal{D}(x),y_{\text{style}})$ $\propto\nabla_{z_{t}}(s\cdot(\text{log}p(z_{t}|\mathcal{D}(x),y_{\text{style}})-\text{log}p(z_{t}|\mathcal{D}(x)))+\text{log}p(z_{t}|\mathcal{D}(x)))$ (6)
$=\nabla_{z_{t}}(s\cdot\text{log}\underbrace{p(y_{\text{style}}|z_{t},\mathcal{D}(x))}_{\text{style blending}}+\text{log}\underbrace{p(z_{t}|\mathcal{D}(x))}_{\text{geometry}}),$ (7)
also due to Bayes’ rule. Here, maximizing $p(z_{t}|\mathcal{D}(x))$ ensure that the optimized 3D representation maintains the geometry, while maximizing $p(y_{\text{style}}|z_{t},\mathcal{D}(x))$ guides the blending of a style that suits the geometry.
@@ 3.2 Multi-View ControlNet
As elaborated in the previous section, our pipeline generates style in the second stage by refining the DMTet, converted from NeRF. We note the limitations of MVDream in refining DMTet, stemming from a lack of understanding of the geometry. To address this, we introduce the Multi-View (MV) ControlNet, crucial for generating fine-grained style by understanding coarse-grained geometry.
@@ 3.2.1 Generating Dataset
To train a depth-aware multi-view model, we construct a dataset using the curated 100K high-quality texts from OpenShape [Liu et al.(2023)Liu, Shi, Kuang, Zhu, Li, Han, Cai, Porikli, and Su], filtered with GPT-4 [OpenAI(2023)] for quality assurance. Initially, these texts are refined using BLIP [Li et al.(2022)Li, Li, Xiong, and Hoi] and Azure Cognition Services for 2D image captions and subsequently enriched by GPT-4 to filter out uninformative content. Using MVDream, we generate multi-view images from these texts, matching MVDream’s camera parameters to maintain its prior knowledge. This involves generating four orthogonal views for each text, each with uniformly distributed azimuth angles and elevation. Specifically, we randomly sample the camera parameters in a range of [0.9, 1.1] for distance, [15, 60] for fov, and [0, 30] for elevation. Then, we generate corresponding depth maps from the generated images using MiDAS [Ranftl et al.(2022)Ranftl, Lasinger, Hafner, Schindler, and Koltun]. This process, completed in 1.5 days on four A40 GPUs, involves texts that either explicitly mention ‘3D’ or have ‘3D asset’ appended as a postfix.
@@ 3.2.2 Training MV-ControlNet
Our goal is to train a depth encoder that incorporates multi-view depth conditions into MVDream, leveraging the model’s 3D consistency and generation quality. Following the approach of ControlNet [Zhang et al.(2023)Zhang, Rao, and Agrawala], which involves training an additional encoder atop a frozen diffusion model, we introduce multi-view depths as additional inputs to the pre-trained MVDream. Specifically, a trainable copy of MVDream’s U-Net [Ronneberger et al.(2015)Ronneberger, Fischer, and Brox] encoder, including 3D attentions, is tasked with encoding depths and integrated into the U-Net decoder using zero-initialized convolutions. We train this encoder with the standard diffusion objective [Ho et al.(2020)Ho, Jain, and Abbeel] formalized as follows:
@@
where $\mathcal{E}$ denotes the frozen VAE encoder [Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer], $\mathcal{D}$ is the trainable depth encoder, $\epsilon_{\phi}$ is the frozen MVDream’s U-Net, $y$ is the text prompt, and $c$ is the camera parameter.
Figure 4: We compare the depth-aware MV-ControlNet with the P2P [Hertz et al.(2023)Hertz, Mokady, Tenenbaum, Aberman, Pritch, and Cohen-Or] approach on MVDream, and also against MV-ControlNet variants trained under edge and normal conditions. On the left, source images are displayed alongside their respective conditions. Among these, the depth-conditioned multi-view images display the most visually appealing results.
@@ 4 Experiments
For both qualitative and quantitative assessments, we have created a benchmark dataset comprising a total of 30 pairs of geometry and style prompts. These pairs span across five categories: Animals, Characters, Foods, General, and Objects. Qualitative samples of our pipeline are presented in Fig. 2.
@@ 4.1 Qualitative Comparison of Multi-View Image Generation
Before evaluating 3D generation, we assess MV-ControlNet’s multi-view image generation capabilities by comparing with several baselines as shown in Fig. 4. The baselines include MVDream using P2P [Hertz et al.(2023)Hertz, Mokady, Tenenbaum, Aberman, Pritch, and Cohen-Or], an image editing method, and MV-ControlNet variants trained under the canny edge and normal maps. Regarding MV-ControlNet variants, we observe that canny edge maps result in unrealistic and 3D-inconsistent images, whereas normal maps provide 3D consistency but show degraded text alignment. Conversely, using depth conditioning proves most effective, leading to visually appealing and well-aligned results. This success motivated us to choose depth-conditioned MV-ControlNet as our model.
Figure 5: In (a), we present comparisons with previous pipelines. Hulk’s geometry from Fig. 1, styled as Ironman, reveals that Magic3D and ProlificDreamer often produce texture artifacts, while Fantasia3D and MVDream are prone to color oversaturation. (b) illustrates the results under various input conditions. (c) shows the refinement process using Magic3D, while (d) highlights the superior results achieved with our ControlDreamer.
Figure 6: Human preference study on editability and text alignment in five prompt domains. ControlDreamer consistently ranks highest in user preferences across all domains.
@@ 4.2 Qualitative Comparison of 3D Generation
We compare against several baselines, namely Magic3D [Lin et al.(2023)Lin, Gao, Tang, Takikawa, Zeng, Huang, Kreis, Fidler, Liu, and Lin], Fantasia3D [Chen et al.(2023b)Chen, Chen, Jiao, and Jia], and ProlificDreamer [Wang et al.(2023)Wang, Lu, Wang, Bao, Li, Su, and Zhu]. All these methods, including our ControlDreamer, employ the DMTet [Shen et al.(2021)Shen, Gao, Yin, Liu, and Fidler] algorithm to refine textured meshes. The results of our stylization experiments using these baseline models are illustrated in Fig. 5. With Magic3D and ProlificDreamer, we observe some instances of failure, characterized by only partial texture changes in the geometry. This issue arises from a misalignment between the prior knowledge embedded in Stable Diffusion for the style prompt (e.g., the Ironman is slim) and the geometry from the previous stage. Our MV-ControlNet addresses this misalignment by understanding the depth information of the geometry, thereby enhancing alignment and coherence. While Fantasia3D and MVDream successfully stylize the overall geometry, they tend to produce results with issues like oversaturation or residual traits from the original geometry prompt. We invite the readers to explore additional results on the project page.
4.3 Quantitative Comparison of 3D Generation
4.3.1 Directional CLIP Similarity
To quantitatively compare the alignment between prompt modifications and image changes in ControlDreamer and MVDream, we evaluate the directional CLIP similarity [Gal et al.(2022)Gal, Patashnik, Maron, Bermano, Chechik, and Cohen-Or] between first and second stage 3D models on our benchmark dataset. This score measures how well the change of rendered images align with the text directions within the CLIP space. We use texts with three view-oriented postfixes: ‘a front view’, ‘a side view’, and ‘a back view’, utilizing four views from each azimuth angles (0°, 90°, 180°, 270°) for the 3D samples.
In Table 2, averaged similarities and standard deviations are presented for each domain, across all six prompt candidates and directions. We employ the OpenCLIP ViT-L/14 model [Cherti et al.(2023)Cherti, Beaumont, Wightman, Wortsman, Ilharco, Gordon, Schuhmann, Schmidt, and Jitsev], trained on the LAION-2B dataset, which is different from the text encoder used in ControlDreamer. Our method demonstrates superior performance, excelling in generating text-aligned 3D models across all evaluated domains in our benchmark. Detailed explanations of the used prompts are presented in the Appendix.
@@ 4.3.2 Human Evaluations
We assess preference scores using Amazon Mechanical Turk (AMT) across five domains in Fig. 6. As a result, the proposed method consistently achieves the highest preference scores in all domains, demonstrating its effectiveness. This assessment allows for a comparative ranking of MVDream, Fantasia3D, Magic3D, and ProlificDreamer. We assess each methodology using 30 3D models from each domain, with 200 human evaluators participating. Each evaluator encounters three random samples and repeats the evaluation task three times, resulting in about 20 evaluations per domain. For additional details, the templates for the human evaluation are included in the Appendix.
@@ 4.4.1 Variants of MV-ControlNet
In Fig. 5 (b), we compare the 3D stylization results using MV-ControlNet variants trained under canny edge, normal, and depth conditions. With canny edge inputs, significant artifacts appear in both RGB and normals. When using normal maps, the RGB colors in the 3D models are less photorealistic. However, depth maps consistently produce 3D models without artifacts in both color and normal maps, showcasing their superiority.